# Multilingual Routing in Mixture-of-Experts

### v1

@@ 1 Introduction
Despite remarkable progress in LLM multilingual capabilities enabled by scaling, the highly English-centric pre- and post-training of most models means performance gaps remain large for most languages beyond a select few. In recent years, significant effort has been devoted to interpreting the internal mechanisms that enable multilinguality. This has generally unveiled shared feature spaces in the middle part of the model, which are pivotal to the cross-lingual transfer of model capabilities (Kojima et al., 2024; Wendler et al., 2024; Bandarkar et al., 2025; Wu et al., 2025). However, this work has been limited to dense LLMs, whereas sparse activation patterns in MoE architectures lead to different computational structures whose impact on feature representations remains unexplored. We discuss relevant literature in Section 2
@@
In this work, we investigate multilingual behavior in mixture-of-experts LLMs. To begin with, a data analysis comparing routing across parallel datasets yields numerous coherent findings. Studying Qwen3-30B-A3B (Yang et al., 2025a), Phi-3.5-MoE (Abdin et al., 2024), GPT-OSS-20B OpenAI (2025), and OLMoE (Muennighoff et al., 2025), we find that, despite their sparsity, they adopt similar mechanisms as dense LLMs; leveraging language-agnostic parameters in intermediate model layers—if anything in a clearer, more modular way. In addition, language performance is strongly correlated to its cross-lingual routing alignment to English. We further highlight how multilingual expert specialization impacts router entropy and token-to-token routing similarity. We present detailed from our routing analysis in Section 4.
@@
We build upon this observation by showing that the model’s ability to call upon shared experts is a key driver of multilingual performance. We investigate this via manual interventions into the MoE block’s forward pass to encourage or discourage the activation of specialized experts. We explore steering the routers in different model layers, intervention strengths, and types of experts. In the end, we find that we can improve multilingual task performance when activating experts important for solving that task in English. We experiment with Qwen3, Phi-3.5-MoE, and GPT-OSS—all fully post-trained and state-of-the-art LLMs—and find that these inference-time interventions consistently yield statistically significant improvements on two tasks requiring domain knowledge, MGSM (Shi et al., 2023) and the medicine subset of Global-MMLU (Singh et al., 2025). We describe our intervention methodology in Section 5 and walk through the results in Section 6.
@@
Through this routing data analysis and the resulting intervention experiments, we demonstrate that improved expert-sharing leads to the generalization of such complex capabilities. By demonstrating that simple inference-time interventions yield substantial improvements, our work reveals a vast potential for improving multilingual performance in MoE LLMs. This result motivates the development of other methods that promote cross-lingual expert sharing, such as during training.
@@ 2 Related Work on Multilingual LLMs
Before the massive scaling of decoder-only LLMs, smaller encoder-decoder models were subject to the curse of multilinguality, where adding more languages hurt performance in other languages due to limited representational capacity (Conneau et al., 2020; Pfeiffer et al., 2022). Cross-lingual embedding alignment was commonplace with such models in order to unify feature spaces and facilitate multilingual generalization (Zhou et al., 2016; Schwenk & Douze, 2017; Ouyang et al., 2021; Patra et al., 2023). But with the shift to decoder-only LLMs, this approach became no longer viable. Nonetheless, these LLMs implicitly learn shared feature representations. As noted previously, many works have concluded through different approaches that the middle decoder layers of an LLM contain joint language representations (albeit English-centric), while the first and last layers primarily map language-specific representations to and from this space (Kojima et al., 2024; Wendler et al., 2024; Tang et al., 2024; Alabi et al., 2024). As models have become larger and more trained, this phenomenon has become more evident (Chen et al., 2025) and modular (Bandarkar et al., 2025). Wu et al. (2025) finds that this semantic space extends beyond natural languages, to encompass numbers, computer languages, and different input modalities.
Extending these findings, recent works argue that stronger representational alignment improves multilingual performance (Ravisankar et al., 2025). This has been demonstrated through methods that steer models towards language-shared representations (Mahmoud et al., 2025; Lu et al., 2025) or away from language-specific representations (Lim et al., 2025; Zhao et al., 2025). Prompting methods that encourage the use of English as a pivot language can also boost cross-lingual transfer (Shi et al., 2023; Zhang et al., 2024; Yong et al., 2025). LCM team et al. (2024) introduces an LLM that autoregresses over language-neutral “concept” embeddings instead of subword tokens, which exhibits strong multilingual generalization.
@@ 3 Mixture-of-Experts Preliminaries
MoE LLMs differ from traditional decoder-only transformer architectures by replacing the multi-layer perceptron (MLP) component of each model layer with $E$ MLPs, referred to as “experts”. For each input token, a router (or “gating network”) calculates a set of logits and sends the token embedding to the top-$K$ experts only. The $K$ output hidden states are then aggregated, typically via a weighted sum. MoE models are often trained with an auxiliary load-balancing loss (Shazeer et al., 2017; Fedus et al., 2022) that penalizes uneven expert utilization, introducing some redundancy in expert specialization.
@@ 4.3 Routing Divergence
For each non-English sequence, we use entropy-normalized Jenson-Shannon divergence ($D_{\mathrm{H\text{-}JS}}$) to compare its expert importance distribution to that of its paired English sequence. Routing entropy consistently decreases across model layers and therefore needs to be accounted for when comparing JS-divergence (a symmetric variant of KL-divergence) across layers. We revisit this trend in Section 4.4 and detail our entropy normalization in Appendix A.4. Finally, we average these divergences across all sequences in the corpus to get a metric for routing divergence from English for each layer and language.
@@
Across all layers, language similarity is also correlated with routing similarity. This is very expected in the first and last layers, where token overlap and structural similarity lends itself to shared parameterization. Even so, the impact of language families continues into the middle layers where we find that even here, related language pairs like (Bengali, Assamese), (Farsi, Arabic), or even (Romanized Arabic, MSA) have much lower routing divergence between them than unrelated pairs like (Bengali, French), and (Oriya, Serbian). And while this confounds our analysis relating language ability and routing alignment, we find that language relationships only explain a small part of the trends.
@@
We additionally explore domains instead of languages to compare. We select AlpaCare-MedInstruct (Zhang et al., 2023) to represent the medical domain, GSM8K-Instruct (Cobbe et al., 2021) the mathematical domain, and the English FLoRes split as the baseline  [^1] We note that we cannot pair sequences, so we take $D_{\mathrm{H\text{-}JS}}$ after averaging expert importance across sequences. . For these domains, routing divergence from the generic domain exhibits the opposite pattern: higher divergence exists in the middle layers (more of a $\cap$-shape). However, these patterns are weaker, as domains are less different than languages. Nevertheless, this suggests separation of parameterization between multilinguality and task-specific capabilities, as has been observed in dense models (Bandarkar et al., 2025). We revisit this language-task modularity (Choenni et al., 2024; Bandarkar & Peng, 2025) in Section 5.2, as it is fundamental to our intervention methodology.
@@ 4.4 Routing Entropy and Consistency
We calculate the entropy $H$ of each routing weight distribution and then average across tokens in the datasets. As tokens pass through the model, routers increasingly know which experts to send them to, perhaps because representations become more refined or experts more diverse (Lo et al., 2025). This lowering entropy occurs for English, but is much stronger for non-English languages, as can be seen for OLMoE in Figure 3. For non-English languages in particular, the final layer displays a major drop. While the entropy graphs look quite different for each model (See Appendix A.5), these broader trends are visible for all.
@@
Routing consistency, itself, also correlates with performance, with the lowest resource languages having the highest token-to-token agreement. High-resource languages have lower consistency, but still mostly above that of English across model layers. This can be explained by the English-heavy pretraining using a load-balancing loss, which would lead to a large number of specialized experts for English tokens. In contrast, multilingual tokens rely on fewer experts, resulting in reduced token-to-token variation. In the last layers, the models will consistently send tokens in the same non-English language to the same expert. Both entropy and our consistency metric reinforce the idea that the model has dedicated multilingual generation experts in the last decoder layers, while English tokens have more possible experts to be routed to.
@@ 5.2 Expert Identification
Figure 5: Plot of the Number of Identified Experts per Layer, with $\tau=0.3$ for Qwen3. The red vertical bars delimit the region in which we intervene.
@@ 5.3 Routing Interventions
Or rather $\min(\bm{z})$ if $k\in\mathbb{A}^{-}$. We note that deactivation is a less significant intervention than activation: deactivation removes one of many options, rather than forcing one of the few activations.
@@ 5.4 Intervention Exploration
The models all differed in their MoE configurations (See Appendix 4.2) and magnitudes of $\Delta$-values, requiring model-specific tuning for the expert-selection threshold $\tau$ and the strength of intervention (e.g., hard or soft). For soft interventions, $|\lambda|=0.5$ tended to be the most successful.
6 Intervention Results
@@
Given the lack of improvement from any deactivation schema, we ultimately turn to the more invasive strategy of boosting or force-activating specialized experts and get interesting results:
@@
As discussed, we identify task experts using English in-domain data and the above methodology. Then we intervene to encourage or force the activation of such experts in the middle layers when evaluating on multilingual splits of the benchmark. As displayed in Tables 1 and 2, this intervention method is very consistent in its positive increase for 3 models and 2 evaluation tasks. These gains can be seen across the diverse range of languages, even a bit bigger for lower-resource languages. While the magnitude is modest (1-2 points), this is substantial given the simplicity of the test-time intervention relative to the massive scale and sophistication of the LLMs. It is also statistically significant when considered across languages. The gains for medicine are less pronounced than for math, likely because the dataset for identification is only vaguely related to the evaluation data.
@@
As foreshadowed by the results on deactivation, we find the most important to be location of the intervention. Optimal configurations vary slightly by model, but generally, intervening on a very small numbers of targeted experts leads to improvements. Qwen3 requires the strictest selection threshold $\tau$, while Phi-3.5-MoE works best with minimal interventions (though with 2 experts-per-token, each intervention has much greater impact). For GPT-OSS on MGSM, the effectiveness of hard-activation with $\tau=0.3$ is surprising because here, an expert active on as low as 35% of domain tokens in English would be force-activated on all tokens during evaluation. Despite these differences, the sensitivity to target layers far exceeds that to these tunable parameters for all models, with performance rapidly degrading when steering outside ourselected layers. This strict delimitation of the language-universal middle layers is notable, and also validates the efficacy of the visualizations from the routing analysis in revealing these boundaries. Combining with other interventions, such as activating multilingual experts in the top and bottom layers, tends to nullify the gains from these precise middle layer interventions. In the ends, our experiments show that only by targeting a small number of task-specific experts in these precise, language-universal middle layers can we achieve consistent positive gains in multilingual performance.
@@ 7 Conclusion and Future Work
All our analyses of sparse activation patterns converge on a key finding: the most important multilingual specialization of MoE experts occurs in early and late model layers, with experts in the middle layers serving as language-universal mechanisms for multilingual generalization. Our manual interventions that steer routers to replicate English activation patterns yield consistent multilingual improvements, suggesting a causal relationship between cross-lingual routing alignment and cross-lingual transfer. These findings collectively motivate future work on methods that enhance cross-lingual routing alignment and the sharing of specialized experts. Additionally, the distinct and modular separability of parameters between language-shared and language-specific functions suggests opportunities for architectural or training approaches that exploit this natural division.
@@ Acknowledgments
The authors acknowledge Tanmay Parekh, Sheriff Issaka, and Yunzhi Yao for valuable discussions during this work.
@@ A.2 Routing Divergence Plots for all Models
Figure A.2: Mean entropy-normalized JS-Div per OLMoE layer for 12 non-English languages. We note OLMoE’s poor multilingual capabilities.

### v2

@@ 1 Introduction
Scaling has driven remarkable progress in LLM multilingual capabilities, yet pre- and post-training are often heavily English-centric. As a result, large performance gaps remain beyond a select few languages. In recent years, significant effort has been devoted to interpreting the internal mechanisms that enable multilinguality. This has generally unveiled shared feature spaces in the middle part of the model, which are pivotal to the cross-lingual transfer of model capabilities (Kojima et al., 2024; Wendler et al., 2024; Bandarkar et al., 2025; Wu et al., 2025). However, this work has been limited to dense LLMs, whereas sparse activation in MoE architectures leads to different computational structures whose impact on feature representations remains unexplored.
@@
In this work, we investigate multilingual behavior in mixture-of-experts LLMs. To begin with, a data analysis comparing routing across parallel datasets yields numerous coherent findings. Studying Qwen3-30B-A3B (Yang et al., 2025a), Phi-3.5-MoE (Abdin et al., 2024), GPT-OSS-20B OpenAI (2025), and OLMoE (Muennighoff et al., 2025), we find that, despite their sparsity, they adopt similar mechanisms as dense LLMs; leveraging language-agnostic parameters in intermediate model layers—if anything, in a clearer, more modular way. In addition, language performance is strongly correlated to its cross-lingual routing alignment to English. We further highlight how multilingual expert specialization impacts router entropy and token-to-token routing similarity.
@@
We build upon this observation by showing that the model’s ability to call upon shared experts is a key driver of multilingual performance. We investigate this via manual interventions into the MoE block’s forward pass to encourage or discourage the activation of specialized experts. We explore steering the routers in different model layers, intervention strengths, and types of experts. In the end, we find that we can improve multilingual task performance when activating experts important for solving that task in English. We experiment with Qwen3, Phi-3.5-MoE, and GPT-OSS—all fully post-trained and state-of-the-art LLMs—and find that these inference-time interventions consistently yield statistically significant improvements on two tasks requiring domain knowledge, MGSM (Shi et al., 2023) and the medicine subset of Global-MMLU (Singh et al., 2025). The specific conditions under which our intervention works provides strong validation that our initial interpretability analysis uncovered verifiable MoE mechanisms for processing multilingual text.
@@
Through this routing data analysis and the resulting intervention experiments, we demonstrate that improved expert-sharing leads to the generalization of complex capabilities. By demonstrating that simple inference-time interventions yield substantial improvements, our work reveals a vast potential for improving multilingual performance in MoE LLMs. This result motivates the development of other methods that promote cross-lingual expert sharing, such as during training.
@@ 2 Related Work on Multilingual LLMs
Before the massive scaling of decoder-only LLMs, smaller encoder-decoder models were subject to the curse of multilinguality, where adding more languages hurt performance in other languages due to limited representational capacity (Conneau et al., 2020; Pfeiffer et al., 2022). Cross-lingual embedding alignment was commonplace with such models in order to unify feature spaces and facilitate multilingual generalization (Zhou et al., 2016; Schwenk & Douze, 2017; Ouyang et al., 2021; Patra et al., 2023). But with the shift to decoder-only LLMs, this approach became no longer viable. Nonetheless, these LLMs implicitly learn shared feature representations. As noted previously, many works have concluded through different approaches that the middle decoder layers of an LLM contain joint language representations (albeit English-centric), while the first and last layers primarily map language-specific representations to and from this space (Kojima et al., 2024; Wendler et al., 2024; Tang et al., 2024; Alabi et al., 2024). As model size and training has been scaled, this phenomenon has become more evident (Chen et al., 2025) and modular (Bandarkar et al., 2025). Wu et al. (2025) finds that this semantic space extends beyond natural languages, to encompass numbers, computer languages, and different input modalities.
Recent work finds that stronger middle-layer alignment correlates with improved multilingual performance (Kargaran et al., 2025; Ravisankar et al., 2025). This relationship appears causal: multilingual performance improves when they are explicitly steered toward language-shared representations (Mahmoud et al., 2025; Lu et al., 2025) or away from language-specific representations (Lim et al., 2025; Zhao et al., 2025). Prompting methods that encourage the use of English as a pivot language can also boost cross-lingual transfer (Shi et al., 2023; Zhang et al., 2024; Yong et al., 2025). By design, LCM team et al. (2024) introduces an LLM that autoregresses over language-neutral “concept” embeddings instead of subword tokens, which exhibits strong multilingual generalization.
@@ 3 Mixture-of-Experts Preliminaries
MoE LLMs differ from traditional decoder-only transformer architectures by replacing the multi-layer perceptron (MLP) component of each model layer with $E$ MLPs, referred to as “experts”. For each input token, a router (or “gating network”) calculates a set of logits and sends the token embedding to the top-$K$ experts only. The $K$ output hidden states are then aggregated, typically via a weighted sum. MoE models are often trained with an auxiliary load-balancing loss (Shazeer et al., 2017; Fedus et al., 2022) that penalizes uneven expert utilization, introducing some redundancy in expert specialization. Token-to-expert mappings are established early in pretraining and are often independent of the surrounding context (Xue et al., 2024).
@@ 4.3 Routing Divergence
For each non-English sequence, we use entropy-normalized Jensen-Shannon divergence ($D_{\mathrm{H\text{-}JS}}$) to compare its expert importance distribution to that of its paired English sequence. Routing entropy consistently decreases across model layers and therefore needs to be accounted for when comparing JS-divergence (a symmetric variant of KL-divergence) across layers. We revisit this trend in Section 4.4 and detail our entropy normalization in Appendix A.4. Finally, we average these divergences across all sequences in the corpus to get a metric for routing divergence from English for each layer and language.
@@
Across all layers, language similarity is also correlated with routing similarity. This is very expected in the first and last layers, where token overlap and structural similarity lends itself to shared parameterization. Even so, the impact of language families continues into the middle layers where we find that even here, related language pairs like (Bengali, Assamese) or (Romanized Arabic, MSA) have much lower routing divergence between them than unrelated pairs like (Bengali, French) and (Oriya, Serbian). And while this confounds our analysis relating language ability and routing alignment, we find that language relationships only explain a small part of the trends. Visualizing divergence from Chinese, French, or other high-resource languages (instead of English) yields very similar plots, although the slight attenuation of the trends reflect the English-centricity of the shared representations.
@@
We additionally explore domains instead of languages to compare. We select AlpaCare-MedInstruct (Zhang et al., 2023) to represent the medical domain, GSM8K-Instruct (Cobbe et al., 2021) the mathematical domain, and the English FLoRes split as the baseline . For these domains, routing divergence from the generic domain exhibits the opposite pattern: higher divergence exists in the middle layers (more of a $\cap$-shape). However, these patterns are weaker, as domains are less different than languages. Nevertheless, this suggests separation of parameterization between multilinguality and task-specific capabilities, as has been observed in dense models (Bandarkar et al., 2025). We revisit this language-task modularity (Choenni et al., 2024; Bandarkar & Peng, 2025) in Section 5.2, as it is fundamental to our intervention methodology.
@@ 4.4 Routing Entropy and Consistency
We calculate the entropy $H$ of each routing weight distribution and then average across tokens in the datasets. As tokens pass through the model, routers increasingly know which experts to send them to, perhaps because representations become more refined or experts more diverse (Lo et al., 2025). This lowering entropy occurs for English, but is much more pronounced for non-English languages, as can be seen for OLMoE in Figure 3. For non-English languages in particular, the final layer displays a major drop. While the entropy graphs look quite different for each model (See Appendix A.5), these broader trends are visible for all. A likely explanation is a small set of experts specialized for non-English generation in the final layers. Meanwhile, English has a broader pool of possible experts to decide from, resulting in higher routing entropy.
@@
Routing consistency, itself, also correlates with performance, with the lowest resource languages having the highest token-to-token agreement. High-resource languages have lower consistency, but still mostly above that of English across model layers. This can be explained by the English-heavy pretraining using a load-balancing loss, which would lead to a large number of specialized experts for English tokens. In contrast, multilingual tokens rely on fewer experts, resulting in reduced token-to-token variation. In the last layers, the models will consistently send tokens in the same non-English language to the same expert. These consistency metrics further support the existence of many fewer multilingual experts in the later layers than those available for English.
@@ 5.2 Expert Identification
Figure 5: Plot of the Number of Identified Experts per Layer, with $\tau=0.3$ for Qwen3. The red vertical bars delimit the region in which we intervene. A different view is displayed in Appendix A.8
@@
This finding provides striking empirical support for the ”functional dissociation” of language and thought in LLMs, as proposed by Mahowald et al. (2024). The clear segregation between multilingual experts and task-specific experts suggests that the model’s internal architecture reflects a modular distinction between formal linguistic competence (i.e. in diverse languages) and functional competence (i.e. for a task).
@@ 5.3 Routing Interventions
Or rather $\min(\bm{z})$ if $k\in\mathbb{A}^{-}$. We note that deactivation is a less significant intervention than activation: deactivation removes one of many options, rather than forcing one of the few activations. Either way, the intervened routing is still far from homogenized across languages. Even hard intervention only forces 1 or 2 experts into the top-K selection (where K is typically 4 or 8). The router still produces a full logit distribution for every token and selects most experts as before. The intervention is simply a targeted nudge to ensure the task-relevant experts are strongly considered.
@@ 5.4 Intervention Exploration
Generally speaking, random or even slightly suboptimal interventions to the router lead to performance degradation. This highlights the sensitivity of intervening in heavily-trained MoE routers and the potential to over-/under-weigh particular experts during inference.
@@
6 Intervention Results
Given the lack of improvement from any deactivation schema, we ultimately turn to the more invasive strategy of boosting or force-activating specialized experts, which leads to compelling findings:
@@ 6 Intervention Results
As discussed, we identify task experts using English in-domain data and the above methodology. Then we intervene to encourage or force the activation of such experts in the middle layers when evaluating on multilingual splits of the benchmark. As displayed in Tables 1 and 2, this intervention method is very consistent in its positive increase for 3 models and 2 evaluation tasks. These gains can be seen across the diverse range of languages, even a bit more pronounced for lower-resource languages. While the magnitude is modest (1-2 accuracy points), this is substantial given the simplicity of the test-time intervention. It is also statistically significant when considered across languages. The gains for medicine are less pronounced than for math, likely because the dataset for identification is less directly aligned to the evaluation data.
@@
Hyperparameters
The models all differed in their MoE configurations (See Appendix 4.2) and magnitudes of $\Delta$-values, requiring model-specific tuning for the expert-selection threshold $\tau$ and the strength of intervention (e.g., hard or soft). For soft interventions, $|\lambda|=0.5$ tended to be the most successful, as larger $\lambda$ likely disrupted the weights for aggregation significantly. Generally, intervening on a very small number of targeted experts leads to improvements. Qwen3 requires the strictest selection threshold $\tau$, while Phi-3.5-MoE works best with minimal interventions (though with 2 experts-per-token, each intervention has much greater impact). For GPT-OSS on MGSM, the effectiveness of hard-activation with $\tau=0.3$ is surprising because here, an expert active on as low as 35% of domain tokens in English would be force-activated on all tokens during evaluation.
@@
Despite the above differences across models, the sensitivity to target layers far exceeds that to these tunable hyperparameters for all models. We leveraged our divergence graphs (Figure 2, Appendix A.2) to determine the layers, and using layers slightly outside of this zone led to significant performance degradation rather than improvement. This strict delimitation is notable; it suggests that while middle layers are naturally language-agnostic, the routing differences elsewhere represent meaningful specialization that our intervention merely disrupts. In addition, these results also validate the efficacy of the visualizations from the routing analysis in revealing these boundaries. Combining with other interventions, such as activating multilingual experts in the top and bottom layers, tends to nullify the gains. In the end, our experiments show that only by targeting a small number of task-specific experts in these precise, language-universal middle layers can we achieve consistent positive gains in multilingual performance.
Overall, the baseline failures and number of conditions under which these interventions do not work highlight the difficulty of intervening in heavily-trained MoE routers without hurting performance. As a result, the consistent improvements from our small-scale test-time intervention are notable. The layer-wise conditions under which it happens imply that the inability of such MoE LLMs to activate similar middle-layer task-experts across languages is a limitation for cross-lingual transfer.
@@ 7 Conclusion and Future Work
All our analyses of sparse activation patterns converge on a key finding: the most important multilingual specialization of MoE experts occurs in early and late model layers, with experts in the middle layers serving as language-universal mechanisms for multilingual generalization. The strongest evidence for multilingual specialization exists in the last layers, where we find lower entropy and high token-to-token consistency in non-English languages. When identifying experts specialized for two other domains, math and medicine, we surprisingly find full separation between the sets of experts for those and for multilinguality. Our manual interventions that steer routers to replicate English activation patterns yield consistent multilingual improvements. The specific conditions under which these interventions lead to positive improvement validate our initial analysis and highlight the existence of the model’s learned mechanism to attempt to leverage similar experts in the middle layers, even if not always successful. These results suggest a causal relationship between cross-lingual routing alignment and cross-lingual transfer. These findings collectively motivate future work on methods that enhance cross-lingual routing alignment and the sharing of specialized experts, such as during training. Additionally, the distinct and modular separability of parameters between language-shared and language-specific functions suggests opportunities for architectural or training approaches that exploit this natural division.
@@ Acknowledgments
The authors acknowledge Tanmay Parekh, Sheriff Issaka, and Yunzhi Yao for valuable discussions at UCLA during this work.
@@ A.2 Routing Divergence Plots for all Models
Figure A.2: Mean entropy-normalized JS-Div per OLMoE layer for 12 non-English languages. We note OLMoE’s poor multilingual capabilities. For languages OLMoE does handle reasonably, French and Chinese, the U-shape is still visible.